the sum of two consecutive even integers is 16 less than three times the larger number. find the integers
Answer by macston(5194)

x=smaller integer; x+2=larger integer
x+(x+2)=3(x+2)-16
2x+2=3x+6-16 Subtract 2x from each side.
2=x-10 Add 10 to each side.
12=x ANSWER 1: The smaller integer is 12.
x+2=14 ANSWER 2: The larger integer is 14.
CHECK
(small integer)+(large integer)=3(large integer)-16
12+14=3(14)-16
26=42-16
26=26